A large logistic manufacturing company has 100 employees specializing in different machines. From them, 20 are proficient in machine P , 45 in machine Q , 30 in machine R , 1 in machine P and Q , 6 in machine Q and R , 5 in machine R and P , and just 1 specialist in all three machines. i) Determine how many of them only specialize in machine Q. ii) And how many of them do not specialize in any of the machines.Β 


1
Expert's answer
2022-03-03T04:46:01-0500

we first define the following

N=100, N(P) =20, N(Q) =45, N(R) =30 , N(P and Q)=1, N(Q and R)=6, N(R and P)=5, N(P and Q and R) = 1


Hence we solve part i) and ii) as below


i) The number of those specializing in machine Q only will be

N(Q) - N(P and Q) + N(P and Q and R) - N(Q and R)

= (45 - 1 +1 - 6) = 39 employees


ii)The number of those who do not specialize in any of the machines equals

N - [ N(P) +N(Q) + N(R) - N(P and Q) - N(P and R) - N(Q and R) + N(P and Q and R) ]

= 100 - [20 +45 +30 -1 -5 - 6 +1]

= 100 - 84

= 16 employees